OpenStudy (anonymous):

32 + (8 - 2) · 4 - 6/3

OpenStudy (zale101):

Please Excuse My Dear Aunt Sally :P Perenthesis, exponents, multiplication, Division, addition, and subtraction P E M D A S

OpenStudy (zale101):

first solve the parenthesis

OpenStudy (zale101):

can you do that?

OpenStudy (anonymous):

no. can you do it as an example?

OpenStudy (zale101):

(8-2) is the parenthesis and if you see that in an equation, solve it first

OpenStudy (zale101):

Just follow what i'm telling you

OpenStudy (anonymous):

k

OpenStudy (zale101):

ok :)

OpenStudy (anonymous):

go on..

OpenStudy (zale101):

solve the parenthesis first and i'll go on to the next step

OpenStudy (anonymous):

32 + 6 x 4 - 6/3?

OpenStudy (zale101):

ok, now the next step should be the exponents, but we don't see any exponents so let's just skip that, Ok now let's do the multiplication part, can u do that?

OpenStudy (anonymous):

32+24-6/3

OpenStudy (zale101):

good, now let's start do the divide, wait. There's no division, let's skip that. So move on to addition, do that.

OpenStudy (zale101):

oh my bad, there's division

OpenStudy (anonymous):

56-6/3

OpenStudy (zale101):

do the division and then add

OpenStudy (zale101):

sorry sorry

OpenStudy (anonymous):

haha yea your right

OpenStudy (anonymous):

your fine:)

OpenStudy (zale101):

I'm sure u already know next steps

OpenStudy (zale101):

i listed them above ^

OpenStudy (anonymous):

idk how to do 56-6/3

OpenStudy (zale101):

from where did u got 56?

OpenStudy (zale101):

don't do the addition yet -.-

OpenStudy (zale101):

32 + (8 - 2) · 4 - 6/3 32 + (6) · 4 - 6/3 32 + 24 - 6/3

OpenStudy (zale101):

u were supposed to do that

OpenStudy (anonymous):

ooh i see

OpenStudy (zale101):

Divide

OpenStudy (zale101):

Add

OpenStudy (anonymous):

54

OpenStudy (anonymous):

correct answer?

OpenStudy (zale101):

yup

OpenStudy (zale101):

What grade r u in?

OpenStudy (anonymous):

cause i got 32+24-2 than thats 54... but its not one of my answer choices???

OpenStudy (anonymous):

wow.... the 32 was supposed to be 3 to the second power... haha my b yo

OpenStudy (anonymous):

thanks for your help i got it now:)

OpenStudy (zale101):

It's an order of operations problem

OpenStudy (zale101):

First you deal with the parenthesis 32 + (8 - 2) x 4 - 6/3 = 32 + 6 x 4 - 6/3 Next you deal with the x and / 32 + 6 x 4 - 6/3 = 32 + 24 - 2 Finally, you deal with the addition (+ )and and subtraction (-) in order 32 + 24 - 2 = 56 - 2 = 54

OpenStudy (zale101):

what are the choices?

OpenStudy (anonymous):

yea i know but i was SUPPOSED to put 3 to the second power + (8-2) x 4 - 6/3, vs 32 + 6 x 4 - 6/3

OpenStudy (anonymous):

and i calculated the correct answer 31

OpenStudy (zale101):

? 31, tell me how did u do it?